OpenStudy (anonymous):

Write a quadratic equation having the given numbers as a solutions. -8 and -2

OpenStudy (anonymous):

IF these are the solutions, then these are what x equals. x = -8 x = -2 Solve for 0 now. x + 8 = 0 x + 2 = 0 Now multiply them. (x + 8)(x + 2) = 0 Can you FOIL this?